生命テクノサイエンス特別実験 10 単位 (必修)

Biological science and technology laboratory  生命テクノサイエンスコース教員

【授業目的】研究テーマに基づいて，実験研究を行ない，実験手法を習得すると ともに，研究成果を発表する． 【授業概要】所属する研究室において修士論文の研究テーマに関する実験を行な うことによって，研究能力を高めるとともに，産業や医療などなどに応用で きる技術を習得する．また研究成果をまとめて，発表する．本科目は工業に 関する科目である． 【授業形式】実験 【キーワード】研究実験，成果の発表 【先行科目】 [先行科目] 【関連科目】『生命テクノサイエンス論文輪講』 (0.5)，『生命テクノサイエンス 実務演習』 (0.5) 【履修要件】なし 【履修上の注意】なし 【到達目標】実験方法の習得 【授業計画】 1. 実験 1 2. 実験 2 3. 実験 3 4. 実験 4 5. 実験 5 6. 実験 6 7. 実験 7 8. 実験 8 9. 実験 9 10. 実験 10 11. 実験 11 12. 実験 12 13. 実験 13 14. 実験 14 15. 実験 15 16. 実験 16 【成績評価基準】修士論文を評価する 【教科書】なし【参考書】論文【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216746 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 生物事務室(M703)

Biological science and technology laboratory 10 units (compulsory)

Teacher of course

Target⟩ According to the subject of study, student carriy out the research to obtain

good ability and skill for reseach and for presntation. Outline⟩ Students carrry out each reseach according to the subject of study to

elevate their reaseach ability and skill in the individual laboratory. In addition, they have to present their reseach reuslts for evaluation.

Style⟩ Experiment �������⟩ reseach work, presentation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Practice for undrstanding scientific papers in bioogical

technology”(0.5), “Internship in Biological science and technology”(0.5) Requirement⟩ N/A Notice⟩ N/A Goal⟩ to learn how to perform experiments

Schedule⟩

1. Experiments 1

2. Experiments 2

3. Experiments 3

4. Experiments 4

5. Experiments 5

6. Experiments 6

7. Experiments 7

8. Experiments 8

9. Experiments 9

10. Experiments 10

11. Experiments 11

12. Experiments 12

13. Experiments 13

14. Experiments 14

15. Experiments 15

16. Experiments 16 Evaluation Criteria⟩ Evaluate rsearch papers for the master degree Textbook⟩ N/A Reference⟩ Research papers

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216746

Student⟩ Able to be taken by only specified class(es)

Contact⟩ ⇒ Administration Office(M703)